Expand (x + 2)(3x^2 + 12) + (x - 2).

Found 2 solutions by greenestamps, timofer:
Answer by greenestamps(13209)



(x + 2)(3x^2 + 12) + (x - 2)

In the first part of the expression, multiply each term of the first polynomial by each term of the second.

x(3x^2) = 3x^3
x(12) = 12x
2(3x^2) = 6x^2
2(12) = 24

The product of the two polynomials is the sum of those partial products

3x^3+6x^2+12x+24

When adding the last part of the expression, combine like terms if there are any.

(3x^3+6x^2+12x+24) + (x-2) = 3x^3+6x^2+(12+1)x+(24-2)

ANSWER: 3x^3+6x^2+13x+22


Answer by timofer(106)